# How Does Critical Batch Size Scale in Pre-training?

### v1

@@ Abstract
Training large-scale models under given resources requires careful design of parallelism strategies. In particular, the efficiency notion of critical batch size, concerning the compromise between time and compute, marks the threshold beyond which greater data parallelism leads to diminishing returns. To operationalize it, we propose a measure of CBS and pre-train a series of auto-regressive language models, ranging from 85 million to 1.2 billion parameters, on the C4 dataset. Through extensive hyper-parameter sweeps and careful control on factors such as batch size, momentum, and learning rate along with its scheduling, we systematically investigate the impact of scale on CBS. Then we fit scaling laws with respect to model and data sizes to decouple their effects. Overall, our results demonstrate that CBS scales primarily with data size rather than model size, a finding we justify theoretically through the analysis of infinite-width limits of neural networks and infinite-dimensional least squares regression. Of independent interest, we highlight the importance of common hyper-parameter choices and strategies for studying large-scale pre-training beyond fixed training durations. [^1] Code available at https://github.com/hlzhang109/critical-batch-size.
@@ 1 Introduction
Efficient optimization is critical in pre-training large models (LMs) at scale (McCandlish et al., 2018; Shoeybi et al., 2019; Kaplan et al., 2020). In particular, large-batch training is key to accelerating training, as it enables more efficient parallelism across hardware accelerators (You et al., 2017; Goyal et al., 2018). Specifically, understanding the scaling behavior of the critical batch size (CBS) is essential for optimizing data parallelism, as it defines the point beyond which increasing the batch size no longer results in significant gains in computational efficiency. Below the CBS, approximately linear scaling is achievable—doubling the batch size can proportionally reduce the number of optimization steps required to reach a target loss. However, beyond this threshold, further increases in batch size do not markedly improve gradient approximation quality and would lead to diminishing returns, making it essential to balance computational efficiency with model performance (Shallue et al., 2019; McCandlish et al., 2018). This trade-off presents a challenge for studying pre-training given resource constraints as practitioners are compelled to navigate difficult decisions in balancing compute, data, and training time.
@@
Figure 1: Optimization efficiency in Chinchilla (left) and controlled (middle, right) settings. To study the effect of critical batch size across different model sizes, we track the relative number of steps required to reach the Chinchilla-optimal target validation loss. In the Chinchilla setting (left), we have a fixed data-to-model ratio $D/N=C_{\text{Chin}}$ for each scale and we use the loss of batch size $B_{\text{opt}}=2^{8}$ training on Chinchilla tokens ($D=C_{\text{Chin}}N$) to decide the target loss, observing CBS increases with scale. However, when controlling for either model size (middle) or data size (right), the growth in target losses becomes independent of model size. In the fixed model size scenario, varying proportions of Chinchilla tokens are used to determine the target losses. In the fixed data size setting, we use the loss of batch size $B_{\text{opt}}=2^{8}$ at the same number of Chinchilla tokens to select target losses across different model sizes.
We investigate the scaling laws governing CBS in the context of autoregressive transformer-based language model pre-training (Vaswani, 2017; Radford et al., 2018). Although CBS lacks a formal definition in the literature, establishing one is essential for enabling its study. Even with a formalism, analyzing CBS during pre-training is challenging due to the intertwined effects of increasing both model and data sizes. In contrast to previous works that use fixed FLOPs settings to study the optimal hyper-parameters with respect to the lowest validation loss (DeepSeek-AI et al., 2024; Besiroglu et al., 2024; Porian et al., 2024), we aim to decouple of effects of data size and model size on the growth of CBS.
We build on the Chinchilla compute-optimal framework (Hoffmann et al., 2022a), defining the target validation loss for each model size. We measure critical batch size as a metric, which represents the batch size that results in certain overhead compared to linear scaling: we first set the target validation loss to be the one at the Chinchilla step for each model size using batch size 256 and we measure the number of steps to reach that loss for each different batch size; then we derive the transition points that incur $20\%$ overhead when doubling the batch size. We fit the laws by conducting extensive experiments on models ranging from 85M to 1.2B parameters. Through controlled comparisons, we find that CBS increases in Chinchilla settings where model size $N$ and data size $D$ are jointly scaled up (Figure 1, left) or we scale up $D$ (or training duration thereafter, which we will use interchangeably) while keeping $N$ fixed (Figure 1, middle). However, we find that CBS remains invariant when scaling up $N$ while keeping $D$ fixed (Figure 1, right), suggesting that CBS is likely invariant to model size $N$ but more likely to be dependent on data size $D$.
Our contributions are three-fold. Conceptually, we formalize the notion of critical batch size and examine the independent effects of both model and data size. Through controlled studies, we propose scaling laws that decouple the growth of critical batch size from model and data size, providing a precise characterization of critical batch size based on data size—an aspect underexplored in previous research. Theoretically, focusing on a simple least square regression via mini-batch SGD, we provide a theoretical justification for the scaling up of critical batch size with increasing dataset size; using maximal update parameterization, we justify that the critical batch size would not increase with model width beyond a point. Empirically, we adaptively select optimal hyper-parameters by training small 151M proxy models to understand their impact on optimization. Our controlled experiments provide insights into a range of common optimization configurations, including the hyper-parameters of the Adam optimizer, transformer context length adjustments, and scaling strategies based on width versus depth, among others. Our finding that CBS depends primarily on data size implies that when scaling up data, one can trade compute for efficiency to reduce serial training time through greater data parallelism due to the increase of CBS.
@@ 2 Experimental Design and Empirical Findings
We first describe the experimental settings. Refer Appendix D includes extra details. Throughout this paper, we use the abbreviations ‘M’ for million, ‘B’ for billion, and ‘T’ for trillion.
@@ 2.1 Experimental Settings
Model and training details. We train a series of autoregressive LMs with a context length of 512 in different sizes ranging from 85M, 151M, 302M, 604M to 1.2B (Appendix D Table 2) on C4 (Raffel et al., 2020) with optimizer-specific hyper-parameters reported in Table 3. We adopt the tokenizer of Eleuther AI’s gpt-neox-20b that has a vocabulary of size 50280. We set the micro batch size smaller than the global one and use gradient accumulation to simulate the effects of large global batch sizes. We focus on distributed data parallelism scenarios where communication is frequent, which simplifies the evaluation and abstracts actual wall clock savings into a total number of optimization steps. More details on optimizer configurations and evaluation strategies are included in Appendix D.
Experimental design and outline. To study CBS in Chinchilla settings, we need to consider target loss on a holdout validation set and measure the amount of optimization steps required to reach it. We consider the validation loss of batch size $256$ at step $t_{\textrm{Chin}}=C_{\textrm{Chin}}*N/(\textrm{ctx\_len}*\textrm{bsz})\approx N/(25*\textrm{bsz})$ for each model size $N$ in Table 2 (context length ctx_len set to be 512 for every model) (Appendix D Table 1). When scaling up model size jointly with data size, the above implies that each model size would have a different target loss.
Achieving such a goal through the procedure above is challenging not only due to the combinatorially many hyper-parameter combinations but also the unknown training dynamics of each model size and batch size. In particular, as our focus is on the number of training steps needed to achieve a target validation loss, traditional learning rate decay strategies typically require predefining the total training duration (Loshchilov and Hutter, 2022; Hu et al., 2024; Hägele et al., 2024; Defazio et al., 2024). To address this, we propose using exponential weight averaging (EWA) (Polyak and Juditsky, 1992) to achieve the desired target validation loss, a simple approach that outperforms other popular choices (Figure 2). Below, we outline several key aspects step-by-step to approach the goal:
@@
    Training beyond fixed durations or token amounts, allowing to resume training from checkpoints until the target validation loss is achieved (Section 2.2).
@@ 2.2 Training Beyond Fixed Durations for Reaching Target Validation Loss
Benchmarking learning rate schedulers. In practice, LMs are usually trained using fixed token budgets (Hoffmann et al., 2022a), which can determine the total number of iterations the training would undergo. This training process can be easily decomposed into learning rate warmup and decay phases so that a lower learning rate is kept at the end of training to enable better optimization. However, our goal is to find the optimal performing run under various hyper-parameters and optimization conditions. This implies a nontrivial decision regarding selecting the maximum training duration (Defazio et al., 2024; Hägele et al., 2024). As training beyond fixed durations is particularly favorable in many large-scale pre-training scenarios, we benchmark recently proposed methods like schedule-free optimizer (Defazio et al., 2024), cosine, warmup-stable-decay schedule (WSD) (Hu et al., 2024) (or trapezoidal (Zhai et al., 2022) and our proposed constant+EWA strategy which maintains a running average of model weights $\xi_{t+1}=\tau\cdot\xi_{t}+(1-\tau)\cdot\theta_{t}$ to improve optimization and use $\xi$ for evaluations.
(a) Scheduler Comparison
(b) Scheduler Evaluation Trajectories
Figure 2: Comparing and accounting for training dynamics. (a) Scheduler comparison ablation. (b) Evaluation trajectories with a large batch size of 8192. Throughout, we adopt Constant+EWA since it performs the best for large batch sizes and avoids setting a fixed training duration beforehand for reaching a target loss.
We show that our constant+EWA strategy can match the efficiency of cosine scheduling and WSD, especially for large batch sizes (Figure 2). While prior research has primarily focused on the generalization benefits of EWA (Izmailov et al., 2018), our findings reveal that EWA can trade memory for optimization efficiency. This is particularly useful in scenarios where a target loss must be achieved, but practitioners are uncertain of the exact mixmum data size to set up the learning rate schedule for training.
@@ 2.3 Ablation on Model Context Length
We adopt 512 as the context length of LMs for all of our experiments but it is unclear how it would impact the efficiency of training and whether the existence of CBS would vary when we enlarge the context length.
So we sweep over several larger windows $2^{10},2^{11},2^{12}$ (Figure 3) using 151M models and find that they are qualitatively similar in terms of CBS while the 512 would be slightly worse in the absolute number of steps that required to reach the target loss.
@@
Overall all models in 4 context lengths have very similar relative optimization efficiency across various batch sizes and thus justifies our use of 512 for all the experiments.
Takeaway on model context length: Different context lengths $(2^{9}\sim 2^{12})$ have similar efficiency w.r.t. batch size.
2.4 Ablation on Model Width and Depth
Model sizes can typically be scaled up in two main ways: by increasing the width, which involves enlarging the hidden size of the MLP, or by increasing the depth, which entails adding more layers to the network. As the main result in Figure 1 only involves a single way for scaling up models (Table 2), e.g. training 604M models this way has 2×2$\times$2 × width than the 151M ones. To explore alternative scaling strategies, we investigate how the model behavior changes when we scale the 604M model by increasing the depth by 4×4$\times$4 × instead (detailed configurations in Table 5).
@@ 2.4 Ablation on Model Width and Depth
Our findings may offer practical insights for scaling models under a fixed token budget that is allocated in proportion to model size. This is particularly relevant because scaling model width is often favored over increasing depth, as wider models tend to be more amenable to parallelization (Shoeybi et al., 2019; Touvron et al., 2023).
@@ 3.1 Formal Definition of Critical Batch Size
Recall that we define CBS as the batch size that leads to a $20\%$ overhead compared to linear scaling. We will formalize it in this section. First of all, define $\mathcal{R}(N,D,B)$ as the best loss achievable for a model of size $N$ using a single pass on $D$ tokens with a batch size $B$. This would be obtained by optimally tuning all other parameters of the optimizer, while keeping $N,D,B$ fixed. Below is the formal definition of CBS:
@@
Let’s define $f_{N,D}(B)$ to be the number of steps required to reach $\mathcal{R}_{\text{opt}}(N,D)$ as a function of batch size $B$. Clearly $f_{N,D}(B_{\text{opt}})=D/B_{\text{opt}}$. To define the Critical Batch Size, $B^{*}(N,D)$, we can define a linear scaling curve $f^{*}_{N,D}(B)=D/B$. $f^{*}$ matches $f$ at $B_{\text{opt}}$ and then scales down linearly as batch size goes up.
$B^{*}(N,D)$ is defined as the maximum batch size $B^{\prime}>B_{\text{opt}}(N,D)$ such that
$f_{N,D}(B^{\prime})\leq 1.2f^{*}_{N,D}(B^{\prime}).$
@@
Figure 6: Scaling laws for model size and token size derived from all the plots in Figure 1. (left) Target losses are selected according to chinchilla steps and we fit the power law of critical batch size with respect to model sizes $N$ (in million) as $B^{*}=93.20*N^{0.47}$. Fixed the model size but only scale up the training token (middle): all the results are based on 302M models but different target losses are selected proportionally to data size and then fit a power law with respect to token size $D$ (in billion) is $B^{*}=22.91*D^{0.47}=94.40*(D/C_{\textrm{Chin}})^{0.47}$ which is very close to the scaling law for the Chinchilla setting (left). Training to reach the goal loss is determined by a fixed data size with different model sizes (right): keeping the data size fixed, based on which select target losses for different models leads to a scaling law $B^{*}=621.341*N^{0.087}$ barely invariant to model size.
@@
We begin our first step by fitting a power law of batch size ($B$) to the absolute number of steps ($Y$) to reach the target loss $\log(Y)=\log(a+\frac{b}{B^{\alpha}})$ and then derive the critical batch size. Then we derive the CBS via $B^{*}=\frac{b}{5a}+1.2B_{\text{opt}}$, which is implied by a transition point where the total amount of data under this batch size would incur 20% overhead compared to linear scaling: $D_{\text{total}}=(a+b/B_{\text{opt}}^{\alpha})*1.2B_{\text{opt}}=(a+b/B^{\alpha})*B$, where $\alpha=1$ and $B_{\text{opt}}$ is set to be 256 that is observed in the linear scaling regime.
This is based on the fact that the constant term would be $20\%$ of linear scaling. We report the parameters fitted to the power law relationship between the number of steps $Y$ and batch size $B$ in Appendix Table 6. We adopt the fixed $\alpha=1$ solution, as both strategies yield nearly identical forecasting results.
@@
Overall, we observe an increase in CBS when scaling up in compute-optimal training. Our results suggest that a critical batch size around $2^{9}$ to $2^{11}$ would be helpful to efficiently optimize models below 1B on Chinchilla-optimal amount of tokens to study other empirical problems. However, it is common for the number of tokens trained to scale proportionally with the model’s parameter count. So it is unclear whether the growth of CBS is because of the increase in (1) model size or (2) the data size/training duration, a question we explore in the next section.
@@
Controlled comparison with the same data size. Firstly, we use the token size 3.072B of 151M models $t_{\textrm{Chin}}$ to record the target validation loss for each model size, and train all the 302M, 604M, 1.2B models with a smaller duration again to reach this target loss. To optimize for performance when training on fewer tokens, we also tune the warmup steps accordingly. Figure 1 (right) shows that all the curves behave similarly and we observe no increase in CBS when enlarging the model size. Then we fit a scaling law that leads to $B^{*}=621.341*N^{0.087}$ (Figure 6, right), implying that CBS is relatively invariant across model sizes.
@@
Controlled comparison with the same model size. Moreover, focusing on the 302M models, we conduct additional experiments by selecting target losses at 0.28×0.28$\times$0.28 ×, 0.5×0.5$\times$0.5 ×, 2×2$\times$2 ×, and 4×4$\times$4 × the Chinchilla step for batch size 256 runs. This setup results in two under-training and two over-training configurations. To achieve optimal performance in the over-training scenarios, we increase the warm-up ratio accordingly, while for the under-training cases, we reduce the warm-up ratio proportionally. Results in Figure 1 (middle) show that as we enlarge the number of tokens being trained on, we see an increase of CBS, similar to what we have observed for training large models on chinchilla target loss. This can also be seen in the forecasted CBS curves shown in Figure 6 (middle).
@@
We also plot the results for scaling both $N$ and $D$ (Figure 1, left) and only scaling $D$ (Figure 1, right) together in Figure 7. Through side-by-side comparison, we see that (i) each pair of curves in the same color overlaps to a large degree, showing that models of different sizes trained on the same amount of data share a similar critical batch size; (ii) for the Chinchilla setting (first column in the legend), we have models in different sizes (85M, 151M, 604M, 1.2B) trained on different amounts of tokens, the critical batch size increases through scale. However, when we control for model size and only vary data size (second column in the legend), we observe a similar increase in critical batch size as scale increases. Therefore, we can qualitatively understand that the increase of CBS is likely to be agnostic to model sizes but due to the increase of training duration.
@@
Figure 7: Controlled comparison by training 302M models on varying amounts of tokens and then comparing with other model sizes trained in similar amount of tokens. Models with the same color or positioned in the same row of the legend represent this comparison. For a fixed token count of 3.072B, we measure the target loss at that step for each model size.
@@
Takeaway on scaling laws for critical batch sizes: Based on the scaling laws and controlled comparisons, we conclude that the increase in CBS in Chinchilla-optimal training is more strongly attributed to extended training durations rather than the increase of model size.

### v2

@@ Abstract
Training large-scale models under given resources requires careful design of parallelism strategies. In particular, the efficiency notion of critical batch size (CBS), concerning the compromise between time and compute, marks the threshold beyond which greater data parallelism leads to diminishing returns. To operationalize it, we propose a measure of CBS and pre-train a series of auto-regressive language models, ranging from 85 million to 1.2 billion parameters, on the C4 dataset. Through extensive hyper-parameter sweeps and careful control of factors such as batch size, momentum, and learning rate along with its scheduling, we systematically investigate the impact of scale on CBS. Then we fit scaling laws with respect to model and data sizes to decouple their effects. Overall, our results demonstrate that CBS scales primarily with data size rather than model size, a finding we justify theoretically through the analysis of infinite-width limits of neural networks and infinite-dimensional least squares regression. Of independent interest, we highlight the importance of common hyper-parameter choices and strategies for studying large-scale pre-training beyond fixed training durations. [^1] Code available at https://github.com/hlzhang109/critical-batch-size.
@@ 1 Introduction
Efficient optimization is critical in pre-training large models (LMs) at scale (McCandlish et al., 2018; Shoeybi et al., 2019; Kaplan et al., 2020). In particular, large-batch training is key to accelerating training, as it enables more efficient parallelism across hardware accelerators (You et al., 2017; Goyal et al., 2018). Specifically, understanding the scaling behavior of the critical batch size (CBS) is essential for optimizing data parallelism, as it defines the point beyond which increasing the batch size may result in computational efficiency degradation. Below the CBS, approximately linear scaling is achievable—doubling the batch size can proportionally reduce the number of optimization steps required to reach a target loss. However, beyond this threshold, further increases in batch size would lead to diminishing returns, making it essential to balance computational efficiency with model performance (Shallue et al., 2019; McCandlish et al., 2018). This trade-off presents a challenge for studying pre-training given resource constraints as practitioners are compelled to navigate difficult decisions in balancing compute, data, and training time.
We investigate the scaling laws governing CBS in the context of autoregressive transformer-based language modeling (Vaswani, 2017; Radford et al., 2018). Analyzing CBS in pre-training is challenging due to the absence of a precise formalism relating it to model and data sizes in the literature (Kaplan et al., 2020; McCandlish et al., 2018). Moreover, the interwined effects of scaling model and data sizes proportionally (Hoffmann et al., 2022a) further complicate this analysis. Although previous works study the effects of batch size on optimization performance (DeepSeek-AI et al., 2024; Besiroglu et al., 2024; Porian et al., 2024), two crucial differences are (1) they do not decouple model size and data size; (2) they focus on optimal batch size that reaches the minimum loss instead of critical batch size. We measure critical batch size as a metric, which represents the batch size that results in certain overhead compared to linear scaling: given a certain target validation loss, we measure the number of steps to reach it for different batch sizes; we derive the transition points that incur certain overhead when doubling the batch size; then we fit scaling laws w.r.t. model size and data size to systematically study the scaling of CBS.
@@
Figure 1: Optimization efficiency and scaling of critical batch size in Chinchilla (left) and controlled (middle, right) settings. To study the effect of CBS across different model sizes, we track the relative number of steps required to reach a certain target validation loss. In the Chinchilla setting (left), we keep the data-to-model size ratio $D/N=C_{\text{Chin}}$ constant and observe that CBS increases with scale. However, when controlling for either model size (middle) or data size (right), the growth in target losses becomes mostly dependent on data size rather than model size (Section 3).
1.1 Empirical Takeaways
Conceptually, we formalize the notion of critical batch size and examine the independent effects of both model and data size. We start by scaling up data size in tandem with model size, as suggested in the Chinchilla compute-optimal framework (Hoffmann et al., 2022a). Through controlled studies, we propose scaling laws that decouple the growth of critical batch size from model and data size, leading to the following takeaways — an aspect underexplored in previous research.
[t] Empirical Takeaways : 1. In Chinchilla settings, CBS increases when model size $N$ and data size $D$ (or training duration thereafter, which we will use interchangeably) are jointly scaled up (Figure 1, left). 2. If we scale up training duration $D$ while keeping $N$ fixed (Figure 1, middle), the critical batch size increases to a similar degree. 3. However, we find that CBS remains nearly invariant when scaling up $N$ while keeping $D$ fixed (Figure 1, right), suggesting that CBS weakly depends on model size $N$ but more strongly depends on data size $D$. 4. Our experiments on small 151M proxy models provide insights into a range of common hyper-parameters and optimization configurations, including transformer context length adjustments, and scaling strategies based on width versus depth, among others.
Overall, our empirical finding that CBS scales primarily with data size implies that when scaling up data, one can reduce serial training time through greater data parallelism due to the increase of CBS, without a loss in computational efficiency that can be measured by floating point operations (FLOPs).
1.2 Theoretical Implications
Theoretically, maximal update parameterization suggests that, beyond a certain point, increasing the width of the neural network (while keeping data size fixed) does not further increase the critical batch size. In contrast, by analyzing a simple least-squares regression with mini-batch SGD, we provide a theoretical basis for how the critical batch size continues to scale with increasing data size. Specifically, we introduce two informal theorems here and refer readers to Section 4 for more details.
Theorem 1 (Informal version of Theorem 2).
In infinite width regimes (Yang and Hu, 2021), training dynamics and performance of the networks become effectively independent of the model size. Consequently, the critical batch size remains nearly invariant when scaling up the model size beyond this point, indicating that larger models do not require proportionally larger batch sizes to achieve optimal training efficiency.
Corollary 1 (Informal version of Corollary 2).
Consider mini-batch SGD with $D$ samples in the least square problems under power-law source and capacity conditions. The CBS, which enables mini-batch SGD to achieve the minimal expected excess risk while ensuring the fastest possible serial runtime, is given by $B^{*}(D)=\Theta(D^{c})$, where the exponent $c\geq 0$ is determined by the exponents of the source and capacity conditions. In the regime where the variance error tends to be dominant, we have $c>0$, indicating CBS grows with data size.
@@ 2 Experimental Design and Empirical Findings
We first describe the experimental settings and refer readers to Appendix D for extra details. Throughout this paper, we use the abbreviations ‘M’ for million, ‘B’ for billion, and ‘T’ for trillion.
@@ 2.1 Experimental Settings
Model and training details. We train a series of autoregressive LMs with a context length of 512 in different sizes ranging from 85M, 151M, 302M, 604M to 1.2B (Appendix D Table 2) on C4 (Raffel et al., 2020) using Adam (Kingma, 2014) with optimizer-specific hyper-parameters reported in Table 3. We adopt the tokenizer of Eleuther AI’s gpt-neox-20b that has a vocabulary of size 50280. We use small 151M proxy models to analyze hyper-parameters in most ablation studies. We set the micro batch size smaller than the global one and use gradient accumulation to simulate the effects of large global batch sizes. We focus on fully synchronized distributed data parallelism scenarios where communication is frequent, which simplifies the evaluation and abstracts actual wall clock savings into a total number of optimization steps. More details on optimizer configurations and evaluation strategies are included in Appendix D.
Experimental design and outline. To study CBS in Chinchilla settings, we need to consider target loss on a holdout validation set and measure the number of optimization steps required to reach it. We consider the optimal batch size as $B_{\text{opt}}$ that is in the linear scaling regime that incurs no efficiency overhead as detailed in Appendix C and model architectures in Table 2. We consider the validation loss of an optimal batch size $B_{\text{opt}}=256$ at step $t_{\textrm{Chin}}=C_{\textrm{Chin}}*N/(\textrm{ctx\_len}*B)$ for each model size $N$ and batch size $B$ with context length ctx_len set to be 512.
When scaling up model size jointly with data size, the above implies that each model size would have a different target loss. Achieving such a goal through the procedure above is challenging not only due to the combinatorially many hyper-parameters but also the unknown training dynamics of each model size and batch size. Below, we outline several key aspects to approach the goal:
@@
    As we focus on the number of training steps needed to achieve a target validation loss, learning rate decay strategies typically require predefining the total training duration (Loshchilov and Hutter, 2022; Hu et al., 2024; Hägele et al., 2024; Defazio et al., 2024). To address this, we propose using exponential weight averaging (EWA) (Polyak and Juditsky, 1992) to achieve the desired target validation loss, a simple approach that matches other popular choices (Figure 2). This enables training beyond fixed durations or data size, allowing to resume training from checkpoints until the target validation loss is achieved (Section 2.2).
@@ 2.2 Training Beyond Fixed Durations for Reaching Target Validation Loss
Figure 2: Comparing and accounting for training dynamics. Throughout, we adopt Constant+EWA since it performs the best for large batch sizes and avoids setting a fixed training duration beforehand for reaching a target loss.
Benchmarking learning rate schedulers. In practice, LMs are usually trained using fixed token budgets (Hoffmann et al., 2022a), which can determine the total number of iterations the training would undergo. This training process can be easily decomposed into learning rate warmup and decay phases so that a lower learning rate is kept at the end of training to enable better optimization. However, our goal is to find the optimal performing run under various hyper-parameters and optimization conditions. This implies a non-trivial decision regarding selecting the maximum training duration (Defazio et al., 2024; Hägele et al., 2024). As training beyond fixed durations is particularly favorable in many large-scale pre-training scenarios, we benchmark recently proposed methods like schedule-free optimizer (Defazio et al., 2024), cosine, warmup-stable-decay schedule (WSD) (Hu et al., 2024) (or trapezoidal (Zhai et al., 2022)) and our proposed constant+EWA strategy which maintains a running average of model weights $\xi_{t+1}=\tau\cdot\xi_{t}+(1-\tau)\cdot\theta_{t}$ to improve optimization and use $\xi$ for evaluations. We show that our constant+EWA strategy can match the efficiency of cosine scheduling and WSD, especially for large batch sizes (Figure 2).
Prior research has shown the generalization (Izmailov et al., 2018) and optimization (Karras et al., 2024) benefits of EWA, our findings further reveal that EWA can trade memory for optimization efficiency in LM pre-training, especially in large-batch regimes. This is useful in scenarios where a target loss must be achieved, but practitioners are uncertain of the exact maximum data size to set up the learning rate schedule for training.
@@ 2.3 Ablation on Model Context Length
We adopt 512 as the context length of LMs for all of our experiments but it is unclear how it would impact the efficiency of training and whether the scaling of CBS would vary when we enlarge the context length. So we sweep over several larger windows $2^{10},2^{11},2^{12}$ (Figure 3) Overall all models in four different context lengths have very similar relative optimization efficiency across various batch sizes and thus justifies our use of 512 for all the experiments.
Takeaway on model context length: Different context lengths $(2^{9}\sim 2^{12})$ have similar scaling w.r.t. batch size.
2.4 Ablation on Model Width and Depth
@@
Model sizes can typically be scaled up in two main ways: by increasing the width, which involves enlarging the hidden size of the multilayer perceptron (MLP), or by increasing the depth, which entails adding more layers to the network. As the main result in Figure 1 only involves a single way for scaling up models (Table 2), e.g. 604M model has 2×2$\times$2 × width than the 151M one. To explore alternative scaling strategies, we investigate how the model behavior changes when we scale the 604M model by increasing the depth by 4×4$\times$4 × instead (detailed configurations in Table 5).
@@
Our findings may offer practical insights for scaling models under a fixed token budget that is allocated in proportion to model size. This is particularly relevant because scaling model width is often favored over increasing depth, as wider models tend to be more amenable to parallelization without incurring additional latency overhead (Shoeybi et al., 2019; Touvron et al., 2023; Erdil, 2024).
@@ 3.1 Formal Definition of Critical Batch Size
Recall that CBS is the transition point where increasing the batch size by a factor of $k$, leads to a reduction in the required number of training steps by a factor that is less than $k$. We now define CBS as the batch size that leads to a $20\%$ overhead compared to linear scaling. First of all, define $\mathcal{R}(N,D,B)$ as the best loss achievable for a model of size $N$ using a single pass on $D$ tokens with a batch size $B$. This would be obtained by optimally tuning all other parameters of the optimizer, while keeping $N,D,B$ fixed. Below is the formal definition of CBS:
@@
We begin our first step by fitting a power law of batch size ($B$) to the absolute number of steps ($Y$) to reach the target loss $\log(Y)=\log(a+\frac{b}{B^{\alpha}})$ and then derive the critical batch size. Then we derive the CBS via $B^{*}=\frac{b}{5a}+1.2B_{\text{opt}}$, which is implied by a transition point where the total amount of data under this batch size would incur 20% overhead compared to linear scaling: $D_{\text{total}}=(a+b/B_{\text{opt}}^{\alpha})*1.2B_{\text{opt}}=(a+b/B^{\alpha})*B$ , where $\alpha=1$, $B_{\text{opt}}$ is set to be 256 chosen to lie within the linear scaling regime as suggested in Appendix C. We report the parameters fitted to the power law relationship between the number of steps $Y$ and batch size $B$ in Appendix Table 6. We adopt the fixed $\alpha=1$ solution, as both strategies yield nearly identical forecasting results.
@@
Overall, we observe an increase in CBS when scaling up in compute-optimal training: In Figure 1 (left), we have target losses selected according to chinchilla steps and we fit the power law of critical batch size with respect to model sizes $N$ (in million) as $B^{*}=93.20*N^{0.47}$. Our results suggest that a critical batch size around $2^{9}$ to $2^{11}$ would be helpful to efficiently optimize models below 1B on Chinchilla-optimal amount of tokens to study other empirical problems. However, it is common for the number of tokens trained to scale proportionally with the model’s parameter count. So it is unclear whether the growth of CBS is because of the increase in (1) model size or (2) the data size/training duration, a question we explore in the next subsection.
@@
Controlled comparison with the same data size. Firstly, we use the Chinchilla token size 3.072B of 151M models $t_{\textrm{Chin}}$ to record the target validation loss for each model size and train all the 302M, 604M, 1.2B models with a smaller duration again to reach these target losses. To optimize for performance when training on fewer tokens, we also tune the warmup steps accordingly. Figure 1 (top right) shows that all the curves behave similarly and we observe almost no increase in CBS when enlarging the model size. Moreover, we fit a scaling law with respect to model size thereafter Figure 1 (bottom right): keeping the data size fixed leads to a scaling law $B^{*}=621.341*N^{0.087}$ weakly dependent on model size.
@@
Controlled comparison with the same model size. Moreover, focusing on the 302M models, we conduct additional experiments by selecting target losses at 0.28×0.28$\times$0.28 ×, 0.5×0.5$\times$0.5 ×, 2×2$\times$2 ×, and 4×4$\times$4 × the Chinchilla step for batch size 256 runs. This setup results in two under-training and two over-training configurations. To achieve optimal performance in the over-training scenarios, we increase the warm-up ratio accordingly, while for the under-training cases, we reduce the warm-up ratio proportionally. Results in Figure 1 (middle) show that as we enlarge the number of tokens being trained on, we see an increase of CBS, similar to what we have observed for training large models on chinchilla target loss. This can also be seen in the forecasted CBS curves shown in Figure 1 (middle) which shows that as we enlarge the number of tokens being trained on, we see an increase of CBS, similar to what we have observed in the Chinchilla setting where model and data size are scaled up proportionally.
@@
We also plot the results for scaling both $N$ and $D$ (Figure 1, left) and only scaling $D$ (Figure 1, right) together in Figure 6. In the side-by-side comparison, we observe the following trends: (i) In the Chinchilla setting (indicated by the first column in the legend), models of various sizes (85M, 151M, 604M, 1.2B) trained on different token amounts exhibit an increase in critical batch size as scale grows. (ii) Additionally, each pair of curves with the same color overlaps significantly, indicating that models of different sizes trained on the same token quantity tend to have similar critical batch sizes. (iii) Finally, when model size is held constant and only the data size (second column in the legend) varies, we also observe an increase in critical batch size with scale. Therefore, we can qualitatively understand that the increase of CBS is likely to be agnostic to model sizes but due to the increase in training duration.
@@
Figure 6: Controlled comparison by training 302M models on varying amounts of tokens and then comparing with other model sizes trained in similar amounts of tokens. Models with the same color or positioned in the same row of the legend represent this comparison. For a fixed token count of 3.072B, we measure the target loss at that step for each model size.
@@
Takeaway on scaling laws for critical batch sizes: Based on the scaling laws and controlled comparisons, we conclude that the increase in CBS in Chinchilla-optimal training is more strongly attributed to extended data size or training durations rather than the increase of model size.